# Detection of Correlated Random Vectors

### v1

@@ 4 Lower Bound via Polynomial Decomposition
By Hardy-Ramanujan Formula in (91), $|\mathsf{Par}(m,\leq_{\infty})|$ is sub-exponential in $m$, and the infinite power sum in (90) converge to a finite number for any $\rho^{2}<1$, which we denote by $g(\rho)$. Since the limit function of a power sum is continuous within the region of convergence (which our case is $(-1,1)$), for any sequence $(\rho_{k})_{k}$ such that $\rho_{k}\to 0$ we have that $g(\rho_{k})\to g(0)=1$. Finally, combining with (90) and (25) we conclude that if $\rho^{2}=o(1)$ we have
@@
For the impossibility of strong detection, we recall a well-known fact regarding total-variation distance.
@@
[Tsy04, Lemma 2.6]) For any sequence of measures $(\mu_{n},\nu_{n})_{n\in\mathbb{N}}$ on a measurable space such that $\nu_{n}$ is absolutely continuous w.r.t. $\mu_{n}$, if
$\mathbb{E}_{\mu_{n}}\left[\left(\frac{d\nu_{n}}{d\mu_{n}}\right)^{2}\right]=O(1),$ (95)
@@
By (91) and (90), if $\rho^{2}=1-\Omega(1)$ there exists $\varepsilon\geq 0$ such that $\rho_{k}^{2}\leq 1-\varepsilon$ for all $k\in\mathbb{N}$, and we have
$\mathbb{E}_{{\cal H}_{0}}[{\cal H}(\mathsf{X},\mathsf{Y})]$ $\leq\sum_{m=0}^{\infty}\left|\mathsf{Par}(m,\leq_{\infty})\right|\rho_{k}^{2m}$ (97)
$\leq\sum_{m=0}^{\infty}\left|\mathsf{Par}(m,\leq_{\infty})\right|(1-\varepsilon)^{2m}=O(1).$ (98)
@@
$\mathsf{R}^{\star}=1-d_{\mathsf{TV}}({\mathbb{P}}_{{\cal H}_{0}},{\mathbb{P}}_{{\cal H}_{1}})=1-\Omega(1).$
@@
$\mathbb{E}_{{\cal H}_{0}}\left[{\cal L}(\mathsf{X},\mathsf{Y})^{2}\right]\geq\sum_{m\in\mathbb{N}}\rho^{2m}=\frac{\rho^{2}}{1-\rho^{2}},$ (99)
which is bounded whenever $\rho^{2}=1-\Omega(1)$, and decays to zero whenever $\rho=o(1)$. This shows that the thresholds for the impossibility of weak and strong detection stated in Theorem 2 and Theorem 1 cannot be further improved by a tighter analysis of $\mathbb{E}_{{\cal H}_{0}}[{\cal L}(\mathsf{X},\mathsf{Y})^{2}]$. This implies that the only case in which our lower bound can be improved is if the Cauchy-Schwartz inequality used in (25) is not asymptotically tight.
@@ 5 Partial Correlation in High Dimensions
Our generalized model is formulated again as a binary hypothesis testing problem. Under the null hypothesis ${\cal H}_{0}$, the Gaussian databases $\mathsf{X}\in\mathbb{R}^{n\times d}$ and $\mathsf{Y}\in\mathbb{R}^{n\times d}$ are generated independently at random with $X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n}\sim{\cal N}(0_{d},\mathbf{I}_{d})$. Under the alternative hypothesis ${\cal H}_{1}$, $k$ out of the $n$ elements from the first database $\mathsf{X}$ are correlated with some other $k$ elements in a randomly and uniformly permuted version of $\mathsf{Y}$. Specifically, fix a parameter $k\leq n$, and let ${\cal K}$ be a set of $k$ indices chosen uniformly at random from $\{1,2,\dots,n\}$, and let $\sigma$ be a permutations of $[n]$, chosen uniformly at random as well. Then, for each $i\in\mathcal{K}$ and $1\leq j\leq d$, under ${\cal H}_{1}$, the random vectors $X_{i}(j)$ and $Y_{\sigma(i)}(j)$ are jointly Gaussian with correlation $\rho\neq 0$. To conclude, we deal with the following decision problem,
@@
Lower bounds.
As in Section 3, we start with our lower bounds. We have the following impossibility guarantees for strong and weak detection, respectively.
@@
Consider the detection problem in (100). For any sequence $(\rho,n,k,d)=(\rho_{\ell},n_{\ell},k_{\ell},d_{\ell})_{\ell}$ such that $\rho^{2}<\frac{1}{d}$ and
$\left(\frac{k}{n}\right)^{2}\prod_{i=1}^{k}\frac{1}{1-(d\rho^{2})^{i}}=O(1),$ (106)
@@
$\mathsf{R}^{\star}=1-\Omega(1).$ (107)
@@
Consider the detection problem in (100). For any sequence $(\rho,n,k,d)=(\rho_{\ell},n_{\ell},k_{\ell},d_{\ell})_{\ell}$ such that $\rho^{2}<\frac{1}{d}$ and
$\left(\frac{k}{n}\right)^{2}\prod_{i=1}^{k}\frac{1}{1-(d\rho^{2})^{i}}=o(1),$ (108)
@@
Let us simplify the above results. Recall that for $d=1$ and $k=n$ we obtained that strong detection is impossible if $\rho^{2}<1$. It turns out that if $k=o(n)$, then the detection problem gets statistically harder in the sense that strong detection becomes impossible even if $\rho^{2}<1-\epsilon(n,k)$, where $\epsilon(n,k)\to 0$, as $n\to\infty$. For example, if $k=O(\log n)$, then
and so (106) is satisfied if
for some $\mathsf{C}>0$, and we note that $(k/n)^{\frac{2}{k}}=o(1)$, for $k=O(\log n)$. The discussion above is similar for the weak detection criterion as well.
Upper bounds.
Next, we move to our upper bounds. From the algorithmic point of view, in the high-dimensional case, we propose the following almost straightforward extension of (10),
@@
${\cal Q}_{\rho,d}$ $\triangleq\mathbb{P}_{Q_{XY}^{\otimes d}}\left[\sum_{\ell=1}^{d}{\cal L}_{\mathsf{I}}(A_{\ell},B_{\ell})\geq d\cdot\tau_{\mathsf{count}}\right].$ (114b)
We can then show the following result. Applying Chernoff’s bound on these tails yield that if the threshold $\tau_{\mathsf{count}}$ is such that $\tau_{\mathsf{count}}\in(-d_{\mathsf{KL}}(Q_{XY}||P_{XY}),d_{\mathsf{KL}}(P_{XY}||Q_{XY}))$, then
@@
We mention here that our result holds for any natural $d\geq 1$, while in [EH23] it is assumed that $d\geq d_{0}$, for some fixed $d_{0}\in\mathbb{N}$ (most notably, excluding the interesting $d=1$ case). Also, interestingly, the threshold $\rho^{2}=1-o(n^{-4/d})$ coincides with the threshold for the recovery problem [DCK19], achieved by the exhaustive maximum-likelihood estimator, while the count test is clearly efficient.
@@
The sum test is by no means novel, and was proposed and analyzed for the case where $k=n$ already in [KN22b]. Here, we analyze its performance for any value of $k$, and use a simpler first and second moment analysis compared to [KN22b]. We have the following result.
@@
Thus, if for example $\frac{k}{n}\to\kappa\in(0,1]$, then we see that $\mathsf{R}(\phi_{\mathsf{sum}})\to 0$ if $\kappa^{2}d\rho^{2}=\omega(1)$, strictly better than the guarantees we obtained for the count test. When $\omega(1)=d=O(\log n)$ the count test is successful only if (at best) $\rho^{2}=1-\Omega(1)$, while the sum test allows for vanishing correlations as $\kappa^{2}d\rho^{2}=\omega(1)$. When $d$ is fixed, however, the sum test is completely useless, and the count test is useful. Finally, in the regime where $k=O(n/\sqrt{\log n})$, it can be seen that the count test have better statistical guarantees compared to the sum test.
Finally, we consider the following straightforward generalization of the comparison test in (20). We define,
@@
Again, such a value exists by the definition of the total-variation distance for centered Gaussian random variables (see, e.g., [DMR23, pg. 10]). We then have the following result.
@@
Consider the detection problem in (1), and the comparison test in (118), with $\theta$ given by (21). If $\rho^{2}=\Omega(1)$ and $k=\Theta(n)$, then $\lim_{n\to\infty}\mathsf{R}(\phi_{\mathsf{comp}})<1$, for any $d\geq 1$.
Finally, we mention here that Theorem 9 holds for any natural $d\geq 1$, while the weak detection guarantee (achieved by the sum test) in [KN22b, EH23] hold under the assumption that $d>60\log 2$.

### v3

@@ 4.3 The proof of the lower bound
By Hardy-Ramanujan Formula in (79), $|\mathsf{Par}(m,\leq_{\infty})|$ is sub-exponential in $m$, and the infinite power series in (78) converges to a finite number, for any $\rho^{2}<1$, which we denote by $g(\rho)$. Since the limit function of a power series is continuous within the region of convergence (which in our case is the interval $(-1,1)$), for any sequence $(\rho_{k})_{k}$, such that $\rho_{k}\to 0$, we have $g(\rho_{k})\to g(0)=1$. When combined with (78) and (32), we conclude that if $\rho^{2}=o(1)$ we have,
@@
For the impossibility of strong detection, we recall the following well-known fact regarding the total-variation distance.
@@
[27, Lemma 2.6]) For any sequence of measures $(\mu_{n},\nu_{n})_{n\in\mathbb{N}}$ on a measurable space such that $\nu_{n}$ is absolutely continuous w.r.t. $\mu_{n}$, if
$\mathbb{E}_{\mu_{n}}\left[\left(\frac{\mathrm{d}\nu_{n}}{\mathrm{d}\mu_{n}}\right)^{2}\right]=O(1),$ (83)
@@
Note that Lemma 2 follows from the inequality $1-d_{\mathsf{TV}}(\mu_{n},\nu_{n})\geq\frac{1/2}{1+\chi^{2}(\mu_{n},\nu_{n})}$. We are now in a position to prove Theorem 1.
@@
By (78) and (79), if $\rho^{2}=1-\Omega(1)$, then there exists $\varepsilon\geq 0$ such that $\rho_{k}^{2}\leq 1-\varepsilon$, for all $k\in\mathbb{N}$, and we have,
$\mathbb{E}_{{\cal H}_{0}}[{\cal L}(\mathsf{X},\mathsf{Y})^{2}]$ $\leq\sum_{m=0}^{\infty}\left|\mathsf{Par}(m,\leq_{\infty})\right|\rho_{k}^{2m}$ (85)
$\leq\sum_{m=0}^{\infty}\left|\mathsf{Par}(m,\leq_{\infty})\right|(1-\varepsilon)^{2m}$ (86)
$\leq\sum_{m=0}^{\infty}c\cdot\frac{1}{4\sqrt{3}m}e^{\pi\sqrt{2m/3}}(1-\varepsilon)^{2m}=O(1),$ (87)
@@
$\mathsf{R}^{\star}=1-d_{\mathsf{TV}}({\mathbb{P}}_{{\cal H}_{0}},{\mathbb{P}}_{{\cal H}_{1}})=\Omega(1).$
@@
$\mathbb{E}_{{\cal H}_{0}}\left[{\cal L}(\mathsf{X},\mathsf{Y})^{2}\right]\geq\sum_{m\in\mathbb{N}}\rho^{2m}=\frac{1}{1-\rho^{2}},$ (88)
which is bounded whenever $\rho^{2}=1-\Omega(1)$, and approaches $1$ whenever $\rho=o(1)$. This shows that the thresholds for the impossibility of weak and strong detection stated in Theorems 12 cannot be further improved by a tighter analysis of $\mathbb{E}_{{\cal H}_{0}}[{\cal L}(\mathsf{X},\mathsf{Y})^{2}]$. This implies that the only scenario in which our lower bound can be improved is if the Cauchy-Schwartz inequality used in (32) is not asymptotically tight.
@@ 5 Partial Correlation in High Dimensions
Our generalized model is formulated again as a binary hypothesis testing problem. Under the null hypothesis ${\cal H}_{0}$, the Gaussian databases $\mathsf{X}\in\mathbb{R}^{n\times d}$ and $\mathsf{Y}\in\mathbb{R}^{n\times d}$ are generated independently at random with $X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n}\sim{\cal N}(0_{d},\mathbf{I}_{d})$. Under the alternative hypothesis ${\cal H}_{1}$, $k$ out of the $n$ elements from the first database $\mathsf{X}$ are correlated with some other $k$ elements in a randomly and uniformly permuted version of $\mathsf{Y}$. Specifically, fix a parameter $k\leq n$, and let ${\cal K}$ be a set of $k$ indices chosen uniformly at random from $\{1,2,\dots,n\}$, and let $\sigma$ be a permutations of $[n]$, chosen uniformly at random as well. Then, for each $i\in\mathcal{K}$ and $1\leq j\leq d$, under ${\cal H}_{1}$, the random variables $X_{ij}$ and $Y_{\sigma(i)j}$ are jointly Gaussian with correlation $\rho\neq 0$. To conclude, we deal with the following decision problem,
@@
$\mathsf{R}^{\star}\triangleq\inf_{\phi:\mathbb{R}^{n\times d}\times\mathbb{R}^{n\times d}\to\{0,1\}}\mathsf{R}(\phi).$ (92)
@@
[…]
Figure 3: An illustration of the correlation structure under the alternative hypothesis ${\cal H}_{1}$. Correlated elements are marked with a similar color, where the remaining mutually independent elements are colorless.
Lower bounds. As in Section 3, we start with our lower bounds. We have the following impossibility guarantees for strong and weak detection, respectively.
@@
Consider the detection problem in (LABEL:eqn:decproblem2). For any sequence $(\rho,n,k,d)=(\rho_{\ell},n_{\ell},k_{\ell},d_{\ell})_{\ell}$ such that $\rho^{2}<\frac{1}{d}$ and
$\left(\frac{k}{n}\right)^{2}\left(\prod_{i=1}^{k}\frac{1}{1-(d\rho^{2})^{i}}-1\right)=O(1),$ (93)
@@
$\mathsf{R}^{\star}=\Omega(1).$ (94)
@@
Consider the detection problem in (LABEL:eqn:decproblem2). For any sequence $(\rho,n,k,d)=(\rho_{\ell},n_{\ell},k_{\ell},d_{\ell})_{\ell}$ such that $\rho^{2}<\frac{1}{d}$ and
$\left(\frac{k}{n}\right)^{2}\left(\prod_{i=1}^{k}\frac{1}{1-(d\rho^{2})^{i}}-1\right)=o(1),$ (95)
@@
Let us simplify the above results. Consider first the case where $k=n$, in which the left-hand-side of (94)–(95) above becomes,
$\prod_{i=1}^{n}\frac{1}{1-(d\rho^{2})^{i}}-1.$ (97)
We also observe that
$\prod_{i=1}^{n}\frac{1}{1-(d\rho^{2})^{i}}\leq\prod_{i=1}^{\infty}\frac{1}{1-(d\rho^{2})^{i}}\triangleq g(d\rho^{2}).$ (98)
The function $g$ at the right-hand-side of (98) is the generating function of integer partitions (see, e.g., [29, Chapter 1.3]), and is given by,
$g(x)=\sum_{m=0}^{\infty}\left|\mathsf{Par}(m,\leq_{\infty})\right|x^{m},$ (99)
where $\mathsf{Par}(m,\leq_{\infty})$ is defined in Definition 2. Note that for any $|x|<1$, $g(x)$ is a well-defined finite number. Hence, by (98), if $d\rho^{2}\leq(1-\varepsilon)$, for some constant $\varepsilon>0$, then
$\prod_{i=1}^{n}\frac{1}{1-(d\rho^{2})^{i}}-1\leq g(1-\varepsilon)=O(1).$ (100)
Similarly, if $d\rho^{2}=o(1)$ then
$\prod_{i=1}^{n}\frac{1}{1-(d\rho^{2})^{i}}-1\leq g(o(1))=1+o(1).$ (101)
Subsequently, (100) and (101) imply that strong and weak detection are impossible if $d\rho^{2}\leq(1-\varepsilon)$ and $d\rho^{2}=o(1)$ respectively. These conclusions coincide with the sharp thresholds in [6], in the regime where $d\to\infty$. Now, when $d$ is fixed, it was proved in [6] that strong detection is impossible as long as $\rho^{2}<\rho^{\star}(d)$, for some function $\rho^{\star}(d)$. Accordingly, our result in (100) gives a new threshold for the impossibility of strong detection, suggesting that strong detection is impossible when $\rho^{2}<(1-\varepsilon)d^{-1}$, for some $\varepsilon>0$, even when $d$ is fixed. This threshold, however, is strictly larger than $\rho^{\star}(d)$, for $d<4$, and thus improves upon [6]. For fixed $d\geq 4$, the threshold $\rho^{\star}$ is strictly larger then $d^{-1}$, and our bound does not improves upon the one in [6]. For $k=n$, we only slightly improve existing results; here, our main contribution is our proof technique, which applies for all asymptotic regimes at once, while in [6] a separate analysis for each asymptotic regime is required. For general values of $k$, we show in Section 5.1 that the analysis and techniques in [6] do not extend to the partly correlated regime, as it is unclear how to evaluate the obtained expression for the likelihood’s second moment (see, Proposition 1). This indicates that in addition to the fact that our analysis arguably simpler as compared to [6], our new technique is also essential for analyzing more complex scenarios, which seem quite complicated to analyze using the standard tools.
Upper bounds. Next, we move to our upper bounds. From the algorithmic point of view, in the high-dimensional case, we propose the following almost straightforward extension of (9),
@@
${\cal Q}_{\rho,d}$ $\triangleq\mathbb{P}_{Q_{XY}^{\otimes d}}\left[\sum_{\ell=1}^{d}{\cal L}_{\mathsf{I}}(A_{\ell},B_{\ell})\geq d\cdot\tau_{\mathsf{count}}\right],$ (104b)
where we recall that $Q_{XY}={\cal N}(\mathbf{0},\Sigma_{0})$ and $P_{XY}={\cal N}(\mathbf{0},\Sigma_{\rho})$. We can then show the following result. Applying Chernoff’s bound on these tails yield that if the threshold $\tau_{\mathsf{count}}$ is such that $\tau_{\mathsf{count}}\in(-d_{\mathsf{KL}}(Q_{XY}||P_{XY}),d_{\mathsf{KL}}(P_{XY}||Q_{XY}))$, then
@@
As mentioned right above Theorem 3, the seemingly similar normalized counting procedure in [6], excludes the $d=1$ case, and in fact requires that $d\geq d_{0}$, for some fixed $d_{0}\in\mathbb{N}$. Our result above holds for any natural $d\geq 1$. Also, interestingly, the threshold $\rho^{2}=1-o(n^{-4/d})$ coincides with the threshold for the recovery problem [2], achieved by the exhaustive maximum-likelihood estimator, while the count test is clearly efficient.
@@
Intuitively, $\phi_{\mathsf{sum}}$ sums all $n^{2}$ possible inner products of feature pairs, and compares this sum to the threshold $\frac{kd|\rho|}{2}$, which is simply the average between the expected values of the sum under the null and alternative hypotheses. The idea behind this test is that under the alternative hypothesis, where the databases are correlated, this sum is larger as compared to the null hypothesis, where the databases are uncorrelated. Note that this test is by no means novel, and was proposed and analyzed for the case where $k=n$ already in [4]. Here, we analyze its performance for any value of $k$, and use a simpler first and second moment analysis as compared to [4]. We have the following result.
@@
Thus, if for example $\frac{k}{n}\to\kappa\in(0,1]$, then we see that $\mathsf{R}(\phi_{\mathsf{sum}})\to 0$ if $\kappa^{2}d\rho^{2}=\omega(1)$, strictly better than the guarantees we obtained for the count test. When $\omega(1)=d=O(\log n)$ the count test is successful only if (at best) $\rho^{2}=1-\Omega(1)$, while the sum test allows for vanishing correlations as $\kappa^{2}d\rho^{2}=\omega(1)$. When $d$ is fixed, however, the sum test is completely useless, and the count test is useful. Also, in the regime where $k=O(n/\sqrt{\log n})$, it can be seen that the count test have better statistical guarantees compared to the sum test
Next, we consider weak detection. As mentioned right above Theorem 4, by definition, any test that achieves strong detection achieves weak detection automatically. As so, the count and sum tests above achieve weak detection under the same conditions stated in Theorems 7 and 8. Nonetheless, as for the $d=1$ case, in the regime where $d$ is fixed, we can obtain stronger performance guarantees under the weak detection criterion. Specifically, consider the (straightforward) generalization of the comparison test in (14). We define,
@@
Again, such a value exists by the definition of the total-variation distance for centered Gaussian random variables (see, e.g., [24, pg. 10]). We then have the following result.
@@
Consider the detection problem in (LABEL:eqn:decproblem), and the comparison test in (108), with $\theta$ given by (15). If $\rho^{2}=\Omega(1)$ and $k=\Theta(n)$, then $\lim_{n\to\infty}\mathsf{R}(\phi_{\mathsf{comp}})<1$, for any $d\geq 1$.
Comparing Theorems 7 and 9, we see that when $d$ is fixed, then weak detection is possible if $\rho^{2}$ is of order constant, while for strong detection our test requires the correlation to converge to unity sufficiently fast. On the other hand, if $d\to\infty$, then the count and sum test clearly have better performance guarantees. Finally, we mention here that Theorem 9 holds for any natural $d\geq 1$, while the weak detection guarantee (achieved by the sum test) in [4, 6] holds under the assumption that $d>60\log 2$. For clarity, we emphasize the detection criterion and asymptotic regime where each one the test above are most relevant/operate the best:
  • Count test: strong detection and $d\in\mathbb{N}$.
  • Comparison test: weak detection and $d\in\mathbb{N}$.
  • Sum test: strong detection and $d\to\infty$.
Weak Detection Strong Detection
Asymptotics Possible Impossible Possible Impossible
[…]
Table 2: A summary of our bounds on $\rho^{2}$, for weak and strong detection, as a function of the asymptotic regime. The one marked with * is for $d>1$, where for $d=1$ only $n^{-d/4}$ is defined. The referenced papers contains result matching those presented for the special case $k=n$.
Discussion. In the regime where $k=o(n)$, our results shows that the detection problem gets statistically harder in the sense that strong detection becomes impossible even if $\rho^{2}<(1-\epsilon(n,k))d^{-1}$, where $\epsilon(n,k)\to 0$, as $n\to\infty$. For example, if $k=O(\log n)$, then
and so (93) is satisfied if
for some $\mathsf{C}>0$. We note that
$\left(\frac{k}{n}\right)^{\frac{2}{k}}=\exp\left(-\Omega\left(\frac{n}{\log(n)}\right)\right)=o(1),$ (112)
which proves that the problem is strictly harder. To see that, consider, for example, the one-dimensional case $d=1$. For any $\varepsilon>0$, by (93) and (112) we have that detection is impossible if
$1-n^{-(4+\varepsilon)}<\rho^{2}\leq 1-\exp\left(-\Omega\left(\frac{n}{\log(n)}\right)\right),$ (113)
where such $\rho$ exist for sufficiently large $n$ since $e^{n/\log(n)}$ decays faster than $n^{-\alpha}$, for all $\alpha$. In particular, there exists a sequence $\rho_{n}$ that satisfies the strong detection condition of the count test in Theorem 3 for $k=n$, for which strong detection is theoretically impossible when $k=O(\log n)$. The discussion above is similar for the weak detection criterion as well. When $k=\Theta(n)$, our lower and upper bounds suggest that the detection problem is roughly statistically the same at least asymptotically. Finally, for general values of $k$, our lower and upper bounds are inconclusive on the exact threshold at which a phase transition for weak/strong detection occurs. However, as shown above, the lower bounds in Theorems 5 and 6 show that in certain regimes, the number $k$ of correlated pairs has a non-trivial influence on the statistical hardness of the problem. This could only be proved using our new method for analysing the likelihood’s second moment. We summarize the above results in Table 2, next to previously known bounds in the literature.